# Uncertainty Quantification for Language Models: A Suite of Black-Box, White-Box, LLM Judge, and Ensemble Scorers

### v3

@@ Abstract
Hallucinations are a persistent problem with Large Language Models (LLMs). As these models become increasingly used in high-stakes domains, such as healthcare and finance, the need for effective hallucination detection is crucial. To this end, we outline a versatile framework for zero-resource hallucination detection that practitioners can apply to real-world use cases. To achieve this, we adapt a variety of existing uncertainty quantification (UQ) techniques, including black-box UQ, white-box UQ, and LLM-as-a-Judge, transforming them as necessary into standardized response-level confidence scores ranging from 0 to 1. To enhance flexibility, we propose a tunable ensemble approach that incorporates any combination of the individual confidence scores. This approach enables practitioners to optimize the ensemble for a specific use case for improved performance. To streamline implementation, the full suite of scorers is offered in this paper’s companion Python toolkit, uqlm. To evaluate the performance of the various scorers, we conduct an extensive set of experiments using several LLM question-answering benchmarks. We find that our tunable ensemble typically surpasses its individual components and outperforms existing hallucination detection methods. Our results demonstrate the benefits of customized hallucination detection strategies for improving the accuracy and reliability of LLMs.
@@ 1 Introduction
Large language models (LLMs) are being increasingly used in production-level applications, often in high-stakes domains such as healthcare or finance. Consequently, there is a rising need to monitor these systems for the accuracy and factual correctness of model outputs. In these sensitive use cases, even minor errors can pose serious safety risks while also leading to high financial costs and reputational damage. A particularly concerning risk for LLMs is hallucination, where LLM outputs sound plausible but contain content that is factually incorrect. Many studies have investigated hallucination risk for LLMs (see Huang et al. (2023); Tonmoy et al. (2024); Shorinwa et al. (2024); Huang et al. (2024) for surveys of the literature). Even recent models, such as OpenAI’s GPT-4.5, have been found to hallucinate as often as 37.1% on certain benchmarks (OpenAI, 2025), underscoring the ongoing challenge of ensuring reliability in LLM outputs.
@@
Uncertainty quantification (UQ) techniques can be used for hallucination detection in a zero-resource fashion, meaning they do not require access to a database of source content, ground truth texts, or internet access. These approaches are typically classified as either black-box UQ, white-box UQ, or LLM-as-a-Judge. Black-box UQ methods exploit the stochastic nature of LLMs and measure semantic consistency of multiple responses generated from the same prompt. White-box UQ methods leverage token probabilities associated with the LLM outputs to compute uncertainty or confidence scores. LLM-as-a-Judge methods use one or more LLMs to evaluate the factual correctness of a question-answer concatenation.
@@
In this paper, we outline a versatile framework for generation-time, zero-resource hallucination detection that practitioners can apply to real-world use cases. To achieve this, we adapt a variety of existing black-box UQ, white-box UQ, and LLM-as-a-Judge methods, applying transformations as necessary to obtain standardized confidence scores that range from 0 to 1. For improved customization, we propose a tunable ensemble approach that incorporates any combination of the individual scorers. The ensemble output is a simple weighted average of these individual components, where the weights can be tuned using a user-provided set of graded LLM responses. This approach enables practitioners to optimize the ensemble for a specific use case, leading to more accurate and reliable hallucination detection. Importantly, our ensemble is extensible, meaning practitioners can expand to include new components as research on hallucination detection evolves.
@@
We evaluate the full suite of UQ scorers on responses generated by four LLMs across various question-answer benchmarks, yielding several empirical insights. Most notably, our tunable ensemble generally outperforms its individual components for hallucination detection. However, the performance ranking of individual scorers varies by dataset, underscoring the value of tailoring methods to specific use cases. Among black-box UQ scorers, entailment-style approaches demonstrate superior performance in our comparisons, while all black-box methods show reduced effectiveness when variation in sampled responses is low. Furthermore, gains from sampling additional responses shrink as the number of candidate responses rises, providing practical guidance for deployment. Lastly, a model’s accuracy on a given dataset positively relates to its performance as a judge of other models’ answers on that dataset.
@@
Finally, this paper is complemented by uqlm, our open-source Python package that provides ready-to-use implementations of all uncertainty quantification methods presented and evaluated in this work. [^1] https://github.com/cvs-health/uqlm uqlm enables practitioners to generate responses and obtain response-level confidence scores by providing prompts (i.e., the questions or tasks for the LLM) along with their chosen LLM. Our framework and toolkit provide researchers and developers a model-agnostic, user-friendly way to implement our suite of UQ-based scorers in real-world use cases, enabling more informed decisions around LLM outputs.
@@ 2 Related Work
Cole et al. (2023) propose evaluating similarity between an original response and candidate responses using exact match-based metrics. In particular, they propose two metrics: repetition, which measures the proportion of candidate responses that match the original response, and diversity, which penalizes a higher proportion of unique responses in the set of candidates. These metrics have the disadvantage of penalizing minor phrasing differences even if two responses have the same meaning. Text similarity metrics assess response consistency in a less stringent manner. Manakul et al. (2023) propose using n-gram-based evaluation to evaluate text similarity. Similar metrics such as ROUGE (Lin, 2004), BLEU (Papineni et al., 2002), and METEOR (Banerjee & Lavie, 2005) have also been proposed (Shorinwa et al., 2024). These metrics, while widely adopted, have the disadvantage of being highly sensitive to token sequence orderings and often fail to detect semantic equivalence when two texts have different phrasing. Sentence embedding-based metrics such as cosine similarity (Qurashi et al., 2020), computed using a sentence transformer such as Sentence-Bert (Reimers & Gurevych, 2019), have also been proposed (Shorinwa et al., 2024). These metrics have the advantage of being able to detect semantic similarity in a pair of texts that are phrased differently. In a similar vein, Manakul et al. (2023) propose using BERTScore (Zhang et al., 2020), based on the maximum cosine similarity of contextualized word embeddings between token pairs in two candidate texts.
Natural Language Inference (NLI) models are another popular method for evaluating similarity between an original response and candidate responses. These models classify a pair of texts as either entailment, contradiction, or neutral. Several studies propose using NLI estimates of $1-P(\text{contradiction})$ or $P(\text{entailment})$ between the original response and a set of candidate responses to quantify uncertainty (Chen & Mueller, 2023; Lin et al., 2024). Zhang et al. (2024) follow a similar approach but instead average across sentences and exclude $P(\text{neutral})$ from their calculations. [^2] Averaging across sentences is done to address long-form responses. Jiang et al. (2024) also address long-form hallucination detection but follow a graph-based approach instead. Other studies compute semantic entropy using NLI-based clustering (Kuhn et al., 2023; Kossen et al., 2024; Farquhar et al., 2024). Qiu & Miikkulainen (2024) estimate density in semantic space for candidate responses.
@@
Manakul et al. (2023) consider two scores for quantifying uncertainty with token probabilities: average negative log probability and maximum negative log probability. While these approaches effectively represent a measure of uncertainty, they lack ease of interpretation, are unbounded, and are more useful for ranking than interpreting a standalone score. Fadeeva et al. (2024) consider perplexity, calculated as the exponential of average negative log probability. Similar to average negative log probability, perplexity also has the disadvantage of being unbounded. They also consider response improbability, computed as the complement of the joint token probability of all tokens in the response. Although this metric is bounded and easy to interpret, it penalizes longer token sequences relative to semantically equivalent, shorter token sequences. Another popular metric is entropy, which considers token probabilities over all possible token choices in a pre-defined vocabulary (Malinin & Gales, 2021; Manakul et al., 2023). Malinin & Gales (2021) also consider the geometric mean of token probabilities for a response, which has the advantage of being bounded and easy to interpret. [^3] For additional white-box uncertainty quantification techniques, we refer the reader to Ling et al. (2024); Bakman et al. (2024); Guerreiro et al. (2023); Zhang et al. (2023); Varshney et al. (2023); Luo et al. (2023); Ren et al. (2023); van der Poel et al. (2022); Wang et al. (2023).
@@
For uncertainty quantification, several studies concatenate a question-answer pair and ask an LLM to score or classify the answer’s correctness. Chen & Mueller (2023) propose using an LLM for self-reflection certainty, where the same LLM is used to judge correctness of the response. Specifically, the LLM is asked to score the response as incorrect, uncertain, or correct, which map to scores of 0, 0.5, and 1, respectively. Similarly, Kadavath et al. (2022) ask the same LLM to state $P(Correct)$ given a question-answer concatenation. Xiong et al. (2024) explore several variations of similar prompting strategies for LLM self-evaluation. More complex variations such as multiple choice question answering generation (Manakul et al., 2023), multi-LLM interaction (Cohen et al., 2023), and follow-up questions (Agrawal et al., 2024) have also been proposed.
@@
Chen & Mueller (2023) propose a two-component ensemble for zero-resource hallucination known as BSDetector. The first component, known as observed consistency, computes a weighted average of two comparison scores between an original response and a set of candidate responses, one based on exact match, and another based on NLI-estimated contradiction probabilities. The second component is self-reflection certainty, which uses the same LLM to judge correctness of the response. In their ensemble, response-level confidence scores are computed using a weighted average of observed consistency and self-reflection certainty. Verga et al. (2024) propose using a Panel of LLM evaluators (PoLL) to assess LLM responses. Rather than using a single large LLM as a judge, their approach leverages a panel of smaller LLMs. Their experiments find that PoLL outperforms large LLM judges, having less intra-model bias in the judgments.
@@ 3.1 Problem Statement
$\hat{h}(y_{i};\theta,\tau)=\mathbb{I}(\hat{s}(y_{i};\theta)<\tau),$ (1)
where $\theta$ could include additional responses generated from $x_{i}$ or other parameters. Note that $\hat{h}(\cdot)=1$ implies a hallucination is predicted. We denote the corresponding ground truth value, indicating whether or not the original response $y_{i}$ actually contains a hallucination, as $h(y_{i})$, where $h$ represents a process to ‘grade’ LLM responses:
$h(y_{i})=\begin{cases}1&y_{i}\text{ contains a hallucination}\\ 0&\text{otherwise}.\end{cases}$ (2)
We adapt our scorers from various techniques proposed in the literature. Each scorer outputs response-level confidence scores to be used for hallucination detection. We transform and normalize scorer outputs, if necessary, to ensure each confidence score ranges from 0 to 1 and higher values correspond to greater confidence. [^4] Note that many of the scorers already have support of $[0,1]$ and hence do not require normalization. Below, we provide details of these various scorers.
@@ 3.2 Black-Box UQ Scorers
$EMR(y_{i};\tilde{\mathbf{y}}_{i})=\frac{1}{m}\sum_{j=1}^{m}\mathbb{I}(y_{i}=\tilde{y}_{ij}).$ (3)
@@
Non-contradiction probability (NCP) is a similar, but less stringent approach. NCP, a component of the BSDetector approach proposed by Chen & Mueller (2023), also conducts pairwise comparison between the original response and each candidate response. In particular, an NLI model is used to classify each pair $(y_{i},\tilde{y}_{ij})$ as entailment, neutral, or contradiction and contradiction probabilities are saved. NCP for original response $y_{i}$ is computed as the average NLI-based non-contradiction probability across pairings with all candidate responses:
$NCP(y_{i};\tilde{\mathbf{y}}_{i})=1-\frac{1}{m}\sum_{j=1}^{m}\frac{\eta(y_{i},\tilde{y}_{ij})+\eta(\tilde{y}_{ij},y_{i})}{2}$ (4)
Above, $\eta(y_{i},\tilde{y}_{ij})$ denotes the contradiction probability of $(y_{i},\tilde{y}_{ij})$ estimated by the NLI model. Following Chen & Mueller (2023) and Farquhar et al. (2024), we use microsoft/deberta-large-mnli for our NLI model.
@@
Another approach for measuring text similarity between two texts is BERTScore (Zhang et al., 2020). Let a tokenized text sequence be denoted as $\textbf{t}=\{t_{1},...t_{L}\}$ and the corresponding contextualized word embeddings as $\textbf{E}=\{\textbf{e}_{1},...,\textbf{e}_{L}\}$, where $L$ is the number of tokens in the text. The BERTScore precision and recall scores between two tokenized texts $\textbf{t},\textbf{t}^{\prime}$ are respectively defined as follows:
@@
$BSC(y_{i};\tilde{\mathbf{y}}_{i})=\frac{1}{m}\sum_{j=1}^{m}2\frac{BertP(y_{i},\tilde{y}_{ij})BertR(y_{i},\tilde{y}_{ij})}{BertP(y_{i},\tilde{y}_{ij})+BertR(y_{i},\tilde{y}_{ij})},$ (6)
@@
$NCS(y_{i};\tilde{\mathbf{y}}_{i})=\frac{1}{2m}\sum_{j=1}^{m}\frac{\mathbf{V}(y_{i})\cdot\mathbf{V}(\tilde{y}_{ij})}{\lVert\mathbf{V}(y_{i})\rVert\lVert\mathbf{V}(\tilde{y}_{ij})\rVert}+\frac{1}{2}.$ (7)
@@
Semantic entropy (SE), proposed by Farquhar et al. (2024), exploits variation in multiple responses to compute a measure of response volatility. The SE approach clusters responses by mutual entailment and, like the NCP scorer, relies on an NLI model. However, in contrast to the aforementioned black-box UQ scorers, semantic entropy does not distinguish between an original response and candidate responses. Instead, it computes a single metric value on a list of responses generated from the same prompt. We consider the discrete version of SE, defined as follows:
$SE(y_{i};\tilde{\mathbf{y}}_{i})=-\sum_{C\in\mathcal{C}}P(C|y_{i},\tilde{\mathbf{y}}_{i})\log P(C|y_{i},\tilde{\mathbf{y}}_{i}),$ (8)
@@
$NSN(y_{i};\tilde{\mathbf{y}}_{i})=1-\frac{SE(y_{i};\tilde{\mathbf{y}}_{i})}{\log(m+1)},$ (9)

### v4

@@ Abstract
Hallucinations are a persistent problem with Large Language Models (LLMs). As these models become increasingly used in high-stakes domains, such as healthcare and finance, the need for effective hallucination detection is crucial. To this end, we outline a versatile framework for closed-book hallucination detection that practitioners can apply to real-world use cases. To achieve this, we adapt a variety of existing uncertainty quantification (UQ) techniques, including black-box UQ, white-box UQ, and LLM-as-a-Judge, transforming them as necessary into standardized response-level confidence scores ranging from 0 to 1. To enhance flexibility, we propose a tunable ensemble approach that incorporates any combination of the individual confidence scores. This approach enables practitioners to optimize the ensemble for a specific use case for improved performance. To streamline implementation, the full suite of scorers is offered in this paper’s companion Python toolkit, uqlm. To evaluate the performance of the various scorers, we conduct an extensive set of experiments using several LLM question-answering benchmarks. We find that our tunable ensemble typically surpasses its individual components and outperforms existing hallucination detection methods. Our results demonstrate the benefits of customized hallucination detection strategies for improving the accuracy and reliability of LLMs.
@@ 1 Introduction
Large language models (LLMs) are being increasingly used in production-level applications, often in high-stakes domains such as healthcare or finance. Consequently, there is a rising need to monitor these systems for the accuracy and factual correctness of model outputs. In these sensitive use cases, even minor errors can pose serious safety risks while also leading to high financial costs and reputational damage. A particularly concerning risk for LLMs is hallucination, where LLM outputs sound plausible but contain content that is factually incorrect. Many studies have investigated hallucination risk for LLMs (see huang2023surveyhallucinationlargelanguage; tonmoy2024comprehensivesurveyhallucinationmitigation; shorinwa2024surveyuncertaintyquantificationlarge; huang2024surveyuncertaintyestimationllms for surveys of the literature). Even recent models, such as OpenAI’s GPT-4.5, have been found to hallucinate as often as 37.1% on certain benchmarks (OpenAI), underscoring the ongoing challenge of ensuring reliability in LLM outputs.
@@
Uncertainty quantification (UQ) techniques can be used for hallucination detection in a closed-book setting, meaning they do not require access to a database of source content, ground truth texts, or internet access. These approaches are typically classified as either black-box UQ, white-box UQ, or LLM-as-a-Judge. Black-box UQ methods exploit the stochastic nature of LLMs and measure semantic consistency of multiple responses generated from the same prompt. White-box UQ methods leverage token probabilities associated with the LLM outputs to compute uncertainty or confidence scores. LLM-as-a-Judge methods use one or more LLMs to evaluate the factual correctness of a question-answer concatenation.
@@
In this paper, we outline a versatile framework for generation-time, closed-book hallucination detection that practitioners can apply to real-world use cases. To achieve this, we adapt a variety of existing black-box UQ, white-box UQ, and LLM-as-a-Judge methods, applying transformations as necessary to obtain standardized confidence scores that range from 0 to 1.111We employ UQ-style signals as a means to produce response-level confidence for hallucination detection, rather than pursuing uncertainty quantification as an end in itself. We do not separate aleatoric and epistemic uncertainty. For improved customization, we propose a tunable ensemble approach that incorporates any combination of the individual scorers. The ensemble output is a simple weighted average of these individual components, where the weights can be tuned using a user-provided set of graded LLM responses. This approach enables practitioners to optimize the ensemble for a specific use case, leading to more accurate and reliable hallucination detection. Importantly, our ensemble is extensible, meaning practitioners can expand to include new components as research on hallucination detection evolves.
@@
We evaluate the full suite of UQ scorers on responses generated by four LLMs across various question-answer benchmarks, yielding several empirical insights. Most notably, our tunable ensemble generally outperforms its individual components for hallucination detection. However, the performance ranking of individual scorers varies by dataset, underscoring the value of tailoring methods to specific use cases. Among black-box UQ scorers, entailment-style approaches demonstrate superior performance in our comparisons. Furthermore, gains from sampling additional responses shrink as the number of candidate responses rises, providing practical guidance for deployment. Lastly, a model’s accuracy on a given dataset positively relates to its performance as a judge of other models’ answers on that dataset.
@@
Finally, this paper is complemented by our open-source Python package, uqlm, that provides ready-to-use implementations of all uncertainty quantification methods presented and evaluated in this work. [^2] The uqlm repository can be found at https://github.com/cvs-health/uqlm. For a detailed description of the software, we refer the reader to bouchard2025uqlmpythonpackageuncertainty. uqlm enables practitioners to generate responses and obtain response-level confidence scores by providing prompts (i.e., the questions or tasks for the LLM) along with their chosen LLM. Our framework and toolkit provide researchers and developers a model-agnostic, user-friendly way to implement our suite of UQ-based scorers in real-world use cases, enabling more informed decisions around LLM outputs.
@@ 2 Related Work
cole2023selectivelyansweringambiguousquestions propose evaluating similarity between an original response and candidate responses using exact match-based metrics. In particular, they propose two metrics: repetition, which measures the proportion of candidate responses that match the original response, and diversity, which penalizes a higher proportion of unique responses in the set of candidates. These metrics have the disadvantage of penalizing minor phrasing differences even if two responses have the same meaning. Text similarity metrics assess response consistency in a less stringent manner. manakul2023selfcheckgptzeroresourceblackboxhallucination propose using n-gram-based evaluation to evaluate text similarity. Similar metrics such as ROUGE (lin-2004-rouge), BLEU (10.3115/1073083.1073135), and METEOR (banerjee-lavie-2005-meteor) have also been proposed (shorinwa2024surveyuncertaintyquantificationlarge). These metrics, while widely adopted, have the disadvantage of being highly sensitive to token sequence orderings and often fail to detect semantic equivalence when two texts have different phrasing. Sentence embedding-based metrics such as cosine similarity (9194665), computed using a sentence transformer such as Sentence-Bert (reimers2019sentencebertsentenceembeddingsusing), have also been proposed (shorinwa2024surveyuncertaintyquantificationlarge). These metrics have the advantage of being able to detect semantic similarity in a pair of texts that are phrased differently. In a similar vein, manakul2023selfcheckgptzeroresourceblackboxhallucination propose using BERTScore (zhang2020bertscoreevaluatingtextgeneration), based on the maximum cosine similarity of contextualized word embeddings between token pairs in two candidate texts.
Natural Language Inference (NLI) models are another popular method for evaluating similarity between an original response and candidate responses. These models classify a pair of texts as either entailment, contradiction, or neutral. Several studies propose using NLI estimates of $1-P(\text{contradiction})$ or $P(\text{entailment})$ between the original response and a set of candidate responses to quantify uncertainty (chen2023quantifyinguncertaintyanswerslanguage; lin2024generatingconfidenceuncertaintyquantification). zhang2024luqlongtextuncertaintyquantification follow a similar approach but instead average across sentences and exclude $P(\text{neutral})$ from their calculations. [^3] Averaging across sentences is done to address long-form responses. jiang2024graphbaseduncertaintymetricslongform also address long-form hallucination detection but follow a graph-based approach instead. Other studies compute semantic entropy using NLI-based clustering (kuhn2023semanticuncertaintylinguisticinvariances; kossen2024semanticentropyprobesrobust; Farquhar2024). qiu2024semanticdensityuncertaintyquantification estimate density in semantic space for candidate responses.
@@
manakul2023selfcheckgptzeroresourceblackboxhallucination consider two scores for quantifying uncertainty with token probabilities: average negative log probability and maximum negative log probability. While these approaches effectively represent a measure of uncertainty, they lack ease of interpretation, are unbounded, and are more useful for ranking than interpreting a standalone score. fadeeva2024factcheckingoutputlargelanguage consider perplexity, calculated as the exponential of average negative log probability. Similar to average negative log probability, perplexity also has the disadvantage of being unbounded. They also consider response improbability, computed as the complement of the joint token probability of all tokens in the response. Although this metric is bounded and easy to interpret, it penalizes longer token sequences relative to semantically equivalent, shorter token sequences. Another popular metric is entropy, which considers token probabilities over all possible token choices in a pre-defined vocabulary (malinin2021uncertaintyestimationautoregressivestructured; manakul2023selfcheckgptzeroresourceblackboxhallucination). malinin2021uncertaintyestimationautoregressivestructured also consider the geometric mean of token probabilities for a response, which has the advantage of being bounded and easy to interpret. [^4] For additional white-box uncertainty quantification techniques, we refer the reader to ling2024uncertaintyquantificationincontextlearning; bakman2024marsmeaningawareresponsescoring; guerreiro2023lookingneedlehaystackcomprehensive; zhang2023enhancinguncertaintybasedhallucinationdetection; varshney2023stitchtimesavesnine; luo2023zeroresourcehallucinationpreventionlarge; ren2023selfevaluationimprovesselectivegeneration; vanderpoel2022mutualinformationalleviateshallucinations; wang2023selfconsistencyimproveschainthought.
@@
For uncertainty quantification, several studies concatenate a question-answer pair and ask an LLM to score or classify the answer’s correctness. chen2023quantifyinguncertaintyanswerslanguage propose using an LLM for self-reflection certainty, where the same LLM is used to judge correctness of the response. Specifically, the LLM is asked to score the response as incorrect, uncertain, or correct, which map to scores of 0, 0.5, and 1, respectively. Similarly, kadavath2022languagemodelsmostlyknow ask the same LLM to state $P(Correct)$ given a question-answer concatenation. xiong2024llmsexpressuncertaintyempirical explore several variations of similar prompting strategies for LLM self-evaluation. More complex variations such as multiple choice question answering generation (manakul2023selfcheckgptzeroresourceblackboxhallucination), multi-LLM interaction (cohen2023lmvslmdetecting), and follow-up questions (agrawal2024languagemodelsknowtheyre) have also been proposed.
@@
chen2023quantifyinguncertaintyanswerslanguage propose a two-component ensemble for zero-resource hallucination known as BSDetector. The first component, known as observed consistency, computes a weighted average of two comparison scores between an original response and a set of candidate responses, one based on exact match, and another based on NLI-estimated contradiction probabilities. The second component is self-reflection certainty, which uses the same LLM to judge correctness of the response. In their ensemble, response-level confidence scores are computed using a weighted average of observed consistency and self-reflection certainty. verga2024replacingjudgesjuriesevaluating propose using a Panel of LLM evaluators (PoLL) to assess LLM responses. Rather than using a single large LLM as a judge, their approach leverages a panel of smaller LLMs. Their experiments find that PoLL outperforms large LLM judges, having less intra-model bias in the judgments.
@@ 3.1 Problem Statement
$\hat{h}(y_{i};\cdot,\tau)=\mathbb{I}(\hat{s}(y_{i};\cdot)<\tau),$ (1)
where $\hat{s}$, and consequently $\hat{h}$, are conditioned on context variables including the prompt $x_{i}$ and, when applicable, additional responses generated from $x_{i}$. [^5] For notational simplicity, we use $\cdot$” as a placeholder here and explicitly write out conditioning variables in the scorer definitions that follow. Note that $\hat{h}(\cdot)=1$ implies a hallucination is predicted. We denote the corresponding ground-truth label indicating whether the original response $y_{i}$ actually contains a hallucination as $h(y_{i};\cdot)$. The function $h$ represents a grading process that compares a generated response $y_{i}$ to a correct reference answer $y_{i}^{*}$ for the same prompt $x_{i}$:
$h(y_{i};y_{i}^{*},x_{i})=\begin{cases}1&y_{i}\text{ is incorrect with respect to }y_{i}^{*}\\ 0&\text{otherwise}.\end{cases}$ (2)
When the correct answer $y_{i}^{*}$ is known, direct comparison yields the most accurate hallucination label. In deployment, however, $y_{i}^{*}$ is not available at generation time, so $h$ cannot be used as a real-time classifier. We therefore treat $h$ as an oracle ground truth used only offline for evaluation and, where applicable, tuning. Our framework approximates $h$ with $\hat{h}$ using uncertainty-based signals available at generation time, consistent with our closed-book setting and without access to $y_{i}^{*}$.
We adapt techniques from the uncertainty-quantification literature to compute response-level confidence for generation-time hallucination detection. Acknowledging the distinction between uncertainty quantification and hallucination detection, we study UQ-based confidence scorers for this purpose and do not distinguish between aleatoric and epistemic uncertainty. We transform and normalize outputs, if necessary, so that each score lies in $[0,1]$ with higher values indicating greater confidence. [^6] Many scorers already output values in $[0,1]$ and therefore do not require normalization. Below, we detail these scorers.
@@ 3.2 Black-Box UQ Scorers
$EMR(y_{i};\tilde{\mathbf{y}}_{i},x_{i})=\frac{1}{m}\sum_{j=1}^{m}\mathbb{I}(y_{i}=\tilde{y}_{ij}).$ (3)
@@
Non-contradiction probability (NCP) is a similar, but less stringent approach. NCP, a component of the BSDetector approach proposed by chen2023quantifyinguncertaintyanswerslanguage, also conducts pairwise comparison between the original response and each candidate response. In particular, an NLI model is used to classify each pair $(y_{i},\tilde{y}_{ij})$ as entailment, neutral, or contradiction and contradiction probabilities are saved. NCP for original response $y_{i}$ is computed as the average NLI-based non-contradiction probability across pairings with all candidate responses:
$NCP(y_{i};\tilde{\mathbf{y}}_{i},x_{i})=1-\frac{1}{m}\sum_{j=1}^{m}\frac{\eta(y_{i},\tilde{y}_{ij})+\eta(\tilde{y}_{ij},y_{i})}{2}$ (4)
Above, $\eta(y_{i},\tilde{y}_{ij})$ denotes the contradiction probability of $(y_{i},\tilde{y}_{ij})$ estimated by the NLI model. [^7] We note that NLI is an asymmetric measure. manakul2023selfcheckgptzeroresourceblackboxhallucination propose a one-directional variant of NCP. We follow chen2023quantifyinguncertaintyanswerslanguage in using the bidirectional formulation to allow for more flexibility in detecting contradictions. Following chen2023quantifyinguncertaintyanswerslanguage and Farquhar2024, we use microsoft/deberta-large-mnli for our NLI model.
@@
Another approach for measuring text similarity between two texts is BERTScore (zhang2020bertscoreevaluatingtextgeneration). Let a tokenized text sequence be denoted as $\textbf{t}=\{t_{1},...t_{L}\}$ and the corresponding contextualized word embeddings as $\textbf{E}=\{\textbf{e}_{1},...,\textbf{e}_{L}\}$, where $L$ is the number of tokens in the text. The BERTScore precision and recall scores between two tokenized texts $\textbf{t},\textbf{t}^{\prime}$ are respectively defined as follows:
@@
$BSC(y_{i};\tilde{\mathbf{y}}_{i},x_{i})=\frac{1}{m}\sum_{j=1}^{m}2\frac{BertP(y_{i},\tilde{y}_{ij})BertR(y_{i},\tilde{y}_{ij})}{BertP(y_{i},\tilde{y}_{ij})+BertR(y_{i},\tilde{y}_{ij})},$ (6)
@@
$NCS(y_{i};\tilde{\mathbf{y}}_{i},x_{i})=\frac{1}{2m}\sum_{j=1}^{m}\frac{\mathbf{V}(y_{i})\cdot\mathbf{V}(\tilde{y}_{ij})}{\lVert\mathbf{V}(y_{i})\rVert\lVert\mathbf{V}(\tilde{y}_{ij})\rVert}+\frac{1}{2}.$ (7)
@@
Semantic entropy (SE), proposed by Farquhar2024, exploits variation in multiple responses to compute a measure of response volatility. The SE approach clusters responses by mutual entailment and, like the NCP scorer, relies on an NLI model. However, in contrast to the aforementioned black-box UQ scorers, semantic entropy does not distinguish between an original response and candidate responses. Instead, it computes a single metric value on a list of responses generated from the same prompt. We consider the discrete version of SE, defined as follows:
$SE(y_{i};\tilde{\mathbf{y}}_{i},x_{i})=-\sum_{C\in\mathcal{C}}P(C|y_{i},\tilde{\mathbf{y}}_{i})\log P(C|y_{i},\tilde{\mathbf{y}}_{i}),$ (8)
@@
$NSN(y_{i};\tilde{\mathbf{y}}_{i},x_{i})=1-\frac{SE(y_{i};\tilde{\mathbf{y}}_{i},x_{i})}{\log(m+1)},$ (9)